OpenStudy (anonymous):

solve the equation over the interval [0,360). 5cot^2theta-cot theta=2

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{5cot^2\theta-cot \theta=2\\~\\ 5cot^2\theta-cot \theta-2=0\\~\\ put~~cot \theta=x\\~\\ 5x^2-x-2=0\\~\\ x=\dfrac{1\pm \sqrt{41}}{10}\\~\\ cot\theta=\dfrac{1\pm \sqrt{41}}{10}\\~\\ tan\theta=\dfrac{10}{1\pm \sqrt{41}}\\~\\ \theta=arctan(\dfrac{10}{1\pm \sqrt{41}})\\~\\ }\end{align}\)

OpenStudy (anonymous):

no that makes no sense. how is that in the interval [0,360)?

OpenStudy (mathmath333):

i think u have to use calculator here,

OpenStudy (mathmath333):

click on the approximate forms

OpenStudy (anonymous):

how are you getting x?

OpenStudy (mathmath333):

there u have to use quadratic formula if \(\Large ax^2+bx+c=0\) then \(\Large x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

ok thats where i was confused. thanks for your help